# Never too Cocky to Cooperate: An FIM and RL-based USV-AUV Collaborative System for Underwater Tasks in Extreme Sea Conditions

### v1

@@ IV-D Experiment Results and Analysis
  • Maintaining the USV at a fixed position at original coordinates (0,0).
  • Maintaining the USV at a fixed position at midpoint coordinates (100,100).
@@ V Conclusion
In this study, we developed a USV–AUV collaborative system to enhance the performance of underwater tasks in extreme sea conditions, which consists of two key components: (1) high-precision multi-AUV localization achieved through USV path planning via FIM optimization, and (2) RL-based multi-AUV cooperative task execution. Comprehensive experiments in the underwater data collection task validated the system’s feasibility and superior performance, while demonstrating effective USV-AUV coordination and robust operation in extreme sea conditions. To support further research in this field, we have open-sourced a simulation code demo at: https://github.com/360ZMEM/USV-AUV-colab .
In the future, we aim to explore online learning and adaptive techniques to enable real-world deployment with real-time environmental adaptation for the USV-AUV system in extreme sea conditions. Our goal is to address practical challenges such as the Sim2Real gap observed in real-world applications.

### v2

@@ IV-D Experiment Results and Analysis
  • Maintaining the USV at a fixed position at the original coordinates (0 m,0 m).
  • Maintaining the USV at a fixed position at midpoint coordinates (100 m,100 m).
@@
Moreover, to quantitatively evaluate the performance and robustness of the proposed FIM-based USV path planning module against alternative RL-based planning strategies commonly used in similar underwater cooperative settings, we designed two comparative baselines in which the USV’s trajectory was determined solely by standalone RL policies using the TD3 algorithm. The first variant (TD3-R/S) employed a sparse reward emphasizing the coverage of target sensor nodes, while the second (TD3-R/D) adopted a dense reward derived from the Signal-to-Noise Ratio (SNR) of the acoustic signals received by the USV from the AUVs. To further assess robustness under computational constraints, we analyzed how reducing the number of iterations in the FIM determinant maximization process influences localization accuracy and planning stability. The quantitative comparisons summarized in Table V and Table VI evaluate SDR, EC, ARPT, and SSN, while Fig. 11 illustrates the evolution of AUV positioning errors under different conditions. As shown in Table V, although TD3-R/D slightly outperforms TD3-R/S in SDR and SSN, both RL-based baselines still underperform the proposed FIM-optimized planner, which achieves the highest SDR (10.78 Mbps) and lowest energy consumption (147.81 W). This demonstrates that the explicit incorporation of information-theoretic guidance significantly enhances data transmission efficiency, control smoothness, and overall trajectory stability.
Finally, we conduct a systematic computational complexity analysis. In our implementation, the FIM-based optimization employs the differential evolution algorithm [43] provided by the scipy library [44], with the number of optimization iterations (denoted by the attribute nit in scipy) used as the primary evaluation metric. Table VI reveals that, as the number of FIM optimization iterations increases, system performance consistently improves across all evaluated metrics. Convergence trends emerge beyond 18 iterations, and optimal results are attained at 24 iterations, where the Spatial Data Rate (SDR) and the Signal-to-Noise Score (SSN) reach 10.82 Mbps and 46.3, respectively. Under this setting, in which 24 iterations are executed on a single CPU core with a fixed clock frequency of 3.6 GHz, representative of typical single-core laptop performance around 2021, the average runtime per optimization is 61 ms. This result, obtained under relatively conservative computational conditions, strongly suggests that real-time execution of the FIM optimization is highly feasible on modern multi-core systems. These numerical trends are further corroborated by Fig. 11, which shows that the localization error of the proposed FIM-based planner remains consistently lower and exhibits reduced temporal fluctuations compared to RL-only baselines. Moreover, increasing the number of optimization iterations progressively suppresses high-frequency deviations in the localization trajectory. Collectively, these results confirm that the FIM-guided optimization mechanism provides superior localization precision, energy efficiency, and stability, demonstrating the reliability and adaptability of the proposed USV–AUV collaborative framework. The observed performance gains clearly stem from the mathematical coupling between information-driven trajectory optimization and RL-based decision-making, validating the effectiveness and generalizability of the proposed bi-level information-structured design.
@@ V Conclusion
In this study, we developed a USV–AUV collaborative system designed to enhance underwater task performance under extreme sea conditions. The framework comprises two core components: (1) high-precision multi-AUV localization, achieved through FIM-based USV path planning, and (2) RL-driven multi-AUV cooperative task execution. Extensive experiments on underwater data collection tasks verified the system’s feasibility and demonstrated its superior coordination efficiency, robustness, and adaptability in challenging marine environments. To facilitate further research and practical development in this field, we have open-sourced the simulation framework at https://github.com/360ZMEM/USV-AUV-colab .
In future work, we aim to extend the proposed USV–AUV collaborative framework from high-fidelity simulation to real-world deployment, addressing the Sim-to-Real gap and enabling online learning and adaptive control for operation under extreme sea conditions. The system, integrating an information-theoretic FIM-based USV planner and a physically interpretable RL controller, will be enhanced with LLM-driven adaptive mechanisms that dynamically adjust reward weights and control parameters based on environmental feedback, improving robustness and generalization in uncertain marine environments. On the experimental side, we are assembling two customized BlueROV2-based AUVs and a USV platform with embedded processors (e.g., NVIDIA Jetson, ARM modules) for real-time inference and communication-efficient coordination. A four-stage validation roadmap, which covers pool, harbor, nearshore, and offshore trials, will progressively assess waterproofing, localization accuracy, communication stability, and control performance, ultimately bridging the gap between simulation and scalable ocean deployment.